### v3

# Pretraining Language Models Using Translationese

Abstract
In this paper, we explore the utility of Translationese as synthetic data created using machine translation for pre-training language models (LMs) for low-resource languages (LRLs). Our simple methodology consists of translating large amounts of web-crawled monolingual documents (clean) into the LRLs, followed by filtering the translated documents using tiny LMs trained on small but clean LRL data. Taking the case of Indian languages, we pre-train LMs from scratch with 28M and 85M parameters, and then fine-tune them for 5 downstream natural language understanding (NLU) and 4 generative (NLG) tasks. We observe that pre-training on filtered synthetic data leads to relative performance drops of only 0.87% for NLU and 2.35% for NLG, compared to pre-training on clean data, and this gap further diminishes upon the inclusion of a small amount of clean data. We also study the impact of synthetic data filtering and the choice of source language for synthetic data generation. Furthermore, evaluating continually pre-trained larger models like Gemma-2B and Llama-3-8B in few-shot settings, we observe that using synthetic data is competitive with using clean data. Our findings suggest that synthetic data shows promise for bridging the pre-training gap between English and LRLs.
Pretraining Language Models Using Translationese
1 Introduction
Large language models (LLMs) Brown et al. (2020); Luccioni et al. (2023); Almazrouei et al. (2023); Lin et al. (2022) have been able to perform very well on downstream tasks like MMLU Hendrycks et al. (2021), Big-Bench Srivastava et al. (2022), etc, and have even started to reach human potential in many of these tasks. But this performance has very largely been credited to their scale and the vast amount of data that they have been fed. Most of these language models (LMs) perform well in languages like English where abundant data is available Kudugunta et al. (2023), but a vast majority of languages don’t have comparable data as compared to English. As a consequence, many LLMs, both monolingual and multilingual, involving these languages still show poor performance for various downstream tasks. For example, the largest open source multilingual model BLOOM Luccioni et al. (2023) covers 46 natural languages spanning 9 language families, but the top 5 languages comprise 74.14% of the data. Even for models like mT5 (Xue et al., 2021), the top 10 of 107 languages account for more than 75.48% of the training data. Despite the benefits of multilingualism Dabre et al. (2020), this data skew still means that low-resource languages will underperform.
Figure 1: Comparing NLU performance in English, Hindi, and Gujarati shows that filtering synthetic data and adding 10% clean data improves models, approaching the performance of those trained only on clean web data.
Fortunately synthetic data is an option and previous works such as, but not limited to, back-translation Sennrich et al. (2016a), sequence distillation Kim and Rush (2016), also known as forward translation, etc. have shown that synthetic data obtained using machine translation (MT) can supplement resource scarcity and can significantly enhance model performance Popović et al. (2020); Gala et al. (2023). However, to the best of our knowledge, there has been no work on showing the effectiveness of synthetic data for pre-training LMs. Furthermore, the quality of synthetic data is also important, which many works take for granted. While round-trip-translation Moon et al. (2020) or referenceless neural quality estimation (QE) Rei et al. (2021) are viable, they either involve twice the compute or a reasonably large model not available for most languages, and this might not be optimal to determine the quality of synthetic documents efficiently. We thus consider TinyLMs Eldan and Li (2023) as an efficient alternative, which have been shown to model documents by their fluent paragraph generation capabilities.
In this paper, we focus on Indic languages such as Hindi, Gujarati, and Marathi, and present a comprehensive study of the utility of synthetic monolingual data, also called translationese Gellerstam (1986), obtained using machine translation (MT) for pre-training LMs. We propose a simple framework that involves training tiny language models, henceforth TinyLMs, on original web-crawled data (clean) and then using them to filter synthetic data. We then compare LMs of different scales pre-trained on clean and synthetic data followed by fine-tuning on natural language understanding (NLP) and generation (NLG) downstream tasks, where we observe that, while unfiltered synthetic data based LMs are inferior compared to LMs trained on clean data, filtering leads to performance comparable to the latter. We further show that tuning these synthetic data LMs on small clean data leads to further improvements. We also show that these trends hold when continually pre-training LLMs such as Gemma-2B and Llama-3-8B.
Figure 2: Overview of our approach to pre-train language models using translationese data. We leverage rich monolingual corpora in the src language and scarce corpora in the tgt language. Our method involves employing a pre-trained machine translation model to translate src to tgt. We then filter, using perplexity, the resulting text using a TinyLM trained solely on clean tgt monolingual data. The filtered synthetic data can be used to further pretrain larger language models.
Our contributions are:
a. A simple framework involving high-quality MT models and TinyLMs trained on clean web-crawled data to mass-produce and filter synthetic data for LM training.
b. Demonstrating the efficacy of language models (up to Llama-3-8B) trained on filtered synthetic data across a range of NLU and NLG tasks for low resource Indic languages.
c. A new document-level monolingual corpora (IndicMonoDoc) consisting of 39.5B tokens worth of monolingual clean document-level data spanning 22 scheduled languages and English [^1] Our code and datasets are released at https://github.com/meetdoshi90/TranslationesePretraining.
2 Related Work
This paper focuses on creating, filtering, and utilizing synthetic data to train TinyLMs.
Monolingual Data: Previous efforts to collect monolingual corpora for Indic languages include the EMILLE/CIIL corpus McEnery et al. (2000), HindMonoCorp Bojar et al. (2014), Leipzig corpus Goldhahn et al. (2012), IndicCorpv1 Kakwani et al. (2020a), and IndicCorpv2 Doddapaneni et al. (2023). While IndicCorpv2 is large, it is sentence-level and suitable for NLU models but not for longer contexts needed by language models. We extend these corpora and demonstrate the effectiveness of using synthetic data.
Synthetic Data Generation and Quality Estimation: Synthetic data aids NLP tasks like back translation for machine translation Sennrich et al. (2016a); Edunov et al. (2018); Marie et al. (2020); Bogoychev and Sennrich (2019); Ni et al. (2022) and native language identification Goldin et al. (2018). However, using synthetic data for pretraining LMs is less explored due to hallucination Maynez et al. (2020) and ungrounded text Thorne et al. (2018). Evaluation methods like RTT BLEU scores are computationally intensive, while others like BARTScore Yuan et al. (2021), T5Score Qin et al. (2023), MQM, and COMET Rei et al. (2020) require large-scale models or human annotations, limiting scalability. Approaches like KenLM Heafield (2011) have been used to filter monolingual corpora based on perplexity.
Transfer Learning and Cross-Lingual Fine-Tuning: Approaches like translate-train, as described by Conneau et al. (2018), involve fine-tuning a multilingual PLM using machine-translated training data and evaluating in the target language. Oh et al. (2022) combined translate-train and translate-test for improved cross-lingual fine-tuning. In contrast, our work focuses on pretraining language models and exploring how synthetic text impacts pretraining and various downstream NLU and NLG tasks.
TinyLMs: Small LMs, even with 10M parameters, produce fluent and consistent text Eldan and Li (2023). Challenges like BabyLM Warstadt et al. (2023) focus on improving LMs within a fixed data budget. We take motivation from this and leverage TinyLMs for efficient filtering of synthetic documents.
3 Methodology
In this section, we describe our framework for leveraging synthetic data for LM training. This process consists of collecting monolingual (clean) data from the web for low-resource languages, training TinyLMs with it, translating clean data from a high resource language such as English into low-resource languages, using the aforementioned TinyLMs to filter synthetic data, and then using this filtered data to train LMs for downstream tasks. Our framework is illustrated in Figure 2.
Figure 3: Language-wise corpora size comparison with IndicCorpv2 Doddapaneni et al. (2023): Stacked Bars
3.1 Collecting Clean Monolingual Corpora
Following Doddapaneni et al. (2023); Rae et al. (2022); Team et al. (2022), for all languages of interest, we a. obtain a list of URLs to be crawled via word-level n-grams passed to a search engine, b. after URL deduplication, we crawl all applicable webpages, c. automatically and manually Ortiz Suárez et al. (2019); Abadji et al. (2022) filter out unwanted text like HTML tags and emoticons, d. use language detection-based (LID) filtering using cld3222https://github.com/google/cld3 and IndicLID-FTN model Madhani et al. (2023a) to discard languages not of interest, e. perform document filtering to drop documents containing offensive text using toxic words list provided by Team et al. (2022), f. merge all the filtered corpus with Wikipedia, OSCAR Ortiz Suárez et al. (2019) and some dumps of mC4 Xue et al. (2021) and finally, g. perform deduplication at paragraph level using Murmurhash algorithm [^3] https://pypi.org/project/mmh3/ with a 128-bit unsigned hash for each monolingual split of the corpora.
We crawl data for English, with Indic context, and 22 Indic languages. As a result, we end up with IndicMonoDoc, with 27.5 billion tokens worth of Indic language documents and 12 billion tokens of English documents for a total of 39.5 billion tokens of clean monolingual data. This is larger than the corpora released by Doddapaneni et al. (2023), surpassing it by almost 2 times. We use IndicMonoDoc for all experiments with clean data. Figure 3 gives an overview of the comparison of IndicMonoDoc. Note that, creation of IndicMonoDoc is important since IndicCorpV2 is a sentence-level corpus, and training LMs need a document-level corpus. It is important to note that we paid special attention to the low-resource languages. In this paper we only use data corresponding to English, Hindi, Gujarati and Marathi. We report additional details of IndicMonoDoc in Appendix E.
3.2 Generating Translationese (Synthetic)
We utilize state-of-the-art MT models like IndicTrans2 Gala et al. (2023) to generate translationese data. Using beam search with a beam value of 5, we translate English tokens from the clean corpus to the target languages. Due to token limits in MT models, we split documents using the Moses Sentence Splitter [^4] https://pypi.org/project/mosestokenizer/ for sentence-level translations, then merge them back into documents. We use the 1B En-Indic version [^5] https://huggingface.co/ai4bharat/indictrans2-en-indic-1B of IndicTrans2 to translate 5B English tokens worth of documents from IndicMonoDoc into translationese data for Hindi, Marathi and Gujarati.
3.3 Tiny Language Models (TinyLMs)
TinyLMs are small language models inspired by Eldan and Li (2023). We use the Transformer architecture Vaswani et al. (2017) and train them with clean monolingual documents. RoPE embeddings Su et al. (2023) are used instead of learned positional encodings for handling long documents. Following Chinchilla scaling laws Hoffmann et al. (2022), we use compute-optimal word tokens. Although it is plausible to train a TinyLM on unfiltered translationese data to filter itself, our preliminary experiments revealed that they favor poor-quality data and hence we avoid this route.
3.4 Synthetic Data Filtering
We use these TinyLMs to filter the generated translationese data. We do this by using perplexity as a measure of document quality score. For language models, perplexity quantifies how well a model predicts a sequence of tokens. A lower perplexity indicates a natural document. It is calculated by:
$\operatorname{PPL}(W)=\exp\left\{-\frac{1}{N}\sum_{i}^{N}\log p_{\theta}\left(w_{i}\mid w_{<i}\right)\right\}$
where the negative log-likelihood measures the error of the model’s predictions. While calculating perplexity over a sequence of tokens, $W\in$ $w_{1},w_{2},\ldots,w_{N}$ we skip the first $s$ tokens where $s=10$, $e=1024$ and calculate loss until only the first $e$ tokens of the document. We find setting $e$ to larger values can lead to higher variance in the document scores due to the size of the TinyLM. Following initial analyses, we choose $s$ and $e$ such that we remove the high uncertainty of the language at the start of an unseen document and avoid penalizing longer documents due to the fragility of the extrapolation ability of TinyLM [^6] During experiments we saw that these TinyLMs can only go up to a certain context length before deteriorating in quality.. Note that it is important to choose $e$ such that the language model gives a uniform estimate of perplexity over an already seen sequence of tokens $\in$ $w_{s},w_{s+1},\ldots,w_{e}$. For our experiments, we use the TinyLMs to score all synthetically generated translationese data and calculate a document score using the above method. Following, Laurençon et al. (2022) we do subsampling by thresholding document perplexity scores except Laurençon et al. (2022) did it using Ken-LM Heafield (2011) and we do it using our TinyLM. We keep the threshold value such that we include enough documents to reach the computed optimal token count for pretraining experiments.
4 Experiments
This section outlines the training procedures and datasets for the models described in Section 3. We pre-train decoder only LMs and fine-tune all models from scratch in monolingual and bilingual settings using the causal language modeling (CLM) objective for NLG tasks and a linear classification head for classification tasks. We specify the dataset samples used for pretraining and fine-tuning, and analyze the effects of synthetic corpora on pretraining.
4.1 Pretraining Data Settings
We refer to translated text or translationese as synthetic or syn and original or web-crawled data as clean throughout our experiments. For the pretraining of all base models, we use the following naming convention to denote our training splits for each model:
XX-clean: This is a clean subset sampled randomly from IndicMonoDoc where XX represents the language English (EN), Hindi (HI) or Gujarati (GU).
syn-XX_yy-unfiltered: Denotes synthetic monolingual documents in XX language generated by using yy as a source during translation.
syn-XX_yy-filtered: Filtered synthetic data.
+10%: Refers to extended pretraining on a cleaned subset of IndicMonoDoc with an additional 10% tokens compared to regular training.
BI-XX-YY Prefix: Denotes bilingual models trained using an equal mixture of monolingual corpora in XX and YY languages. We append an _syn prefix to either XX or YY if a synthetic version of that language is employed in training, and a -parallel/nonparallel tag to denote whether a parallel version of XX and YY are used or not.
Note, for each split we only use the number of tokens that are required to reach the point of optimality Hoffmann et al. (2022) by the language model.
4.2 Implementation and Training
Tokenizer: We use a common byte-pair-encoding (BPE) Sennrich et al. (2016b) tokenizer using Sentencepiece [^7] https://github.com/google/sentencepiece for all experiments. We train a shared vocabulary of 56k subwords between three languages, English, Hindi, and Gujarati by using 5 Million randomly sampled sentences per language and upsampling for Gujarati.
TinyLMs: We use Pytorch Lightning [^8] https://lightning.ai/docs/pytorch/stable/ for our implementations and train TinyLMs as described in Section 3.3 for filtering. We use hidden sizes of 768 and have two variants, one with 4 layers (mini) and one with 12 layers (base; same as GPT2-base) with 28M and 85M non-embedding parameters respectively. The mini models are trained on clean data with sequence lengths of 4096999We keep long sequence lengths to be able to handle long documents for filtering. (mini-4k) for filtering synthetic documents as described in Section 3.4. On the other hand, for our main pre-training and downstream fine-tuning experiments, we train mini and base models with sequence lengths of 1024 (mini-1k and base-1k). Following Hoffmann et al. (2022) we use 2.4 billion word tokens per language to compute optimal training of base models. Since Gujarati has only 900M tokens in our dataset, whenever Gujarati is involved as the target, we train only the mini-1k model. For models involving English and Hindi, we train both mini and base models. Additional details are in Appendix B.
4.3 Downstream Tasks and Evaluation
We finetune the mini-1k and base-1k models for classification, regression, and generation tasks. Hyperparameter tuning is performed using the validation set for models trained only with clean data, and this process is repeated for different data splits. More details on hyperparameters and evaluation can be found in Appendix B. Primary scores are reported on IndicGLUE Kakwani et al. (2020a) and IndicXNLI (iXNLI) Aggarwal et al. (2022) for Hindi and Gujarati, and the GLUE benchmark validation set Wang et al. (2018) for English. We also experiment with other generation tasks like CNN-Dailymail Nallapati et al. (2016), DailogSum Chen et al. (2021), XL-Sum Hasan et al. (2021), IndicNLG Kumar et al. (2022), FLoRes-200 Team et al. (2022), IN22-Conv & IN22-Gen Gala et al. (2023) and use standard evaluation metrics suitable for each task like accuracy, f1-score, Rouge-L Lin (2004) and chrF++ Popović (2017). Further details about each of the evaluation datasets can be found in Appendix B.1.
5 Results
We now present our results which help establish the utility of synthetic data for language modeling.
5.1 Main Results
In this section, we present results for Hindi, Gujarati, and English language models trained on clean data, as well as synthetic data generated from translations. We demonstrate the impact of filtering and adding additional clean data for extended pretraining of LMs trained solely on synthetic text. Additionally, we observe the effect of using the clean source text along with its translations (synthetic parallel documents) on downstream tasks. We follow the naming convention for different data splits as specified in Section 4. We provide details for the pretraining of each model in Appendix B. We provide additional results in Appendix A.
Filtered Synthetic Data is Competitive with Web Scraped Data: The results in Table LABEL:tab:hi-gu-main and 2 indicate that syn-HI_en-unfiltered, syn-GU_en-unfiltered, and syn-EN_hi-unfiltered exhibit lower downstream performance compared to their filtered counterparts: syn-HI_en-filtered, syn-GU_en-filtered, and syn-EN_hi-filtered, respectively. It is evident that filtering the synthetic documents using TinyLMs significantly improves the performance of both NLU and NLG tasks. We also observe that for tasks like CoLA Warstadt et al. (2019), language models trained solely on synthetic data lag behind when compared to other tasks as seen in Table 8 of Appendix A. This suggests that synthetic corpora may lack certain important elements necessary for language models to perform competitively in linguistic acceptability tasks, as opposed to LMs trained on clean, non-synthetic corpora. Results for base-1k for English are presented in Table 8 in Appendix A because we focus our attention on Indic languages.
sst2 cola mrpc qnli qqp rte mnli-m mnli-mm stsb
Model acc mcc f1 acc f1 acc acc acc pearson Avg.
Original EN-clean 87.95 25.59 83.84 78.83 80.78 64.62 71.6 71.69 73.48 70.93
syn-EN_hi-unfiltered 87.53 19.77 79.02 76.49 77.96 55.4 69.65 70.14 67.37 67.04
syn-EN_hi-filtered 87.61 22.81 81.95 77.63 80.57 56.31 70.19 70.89 69.29 68.58
[…]
syn-EN_gu-unfiltered 83.11 17.66 78.53 66.01 77.68 53.6 63.21 64.55 27.33 59.08
syn-EN_gu-filtered 85.66 21.15 81.45 66.35 77.36 54.15 66.27 65.72 26.16 60.47
[…]
Table 2: Effect of source selection for generating synthetic data on the dev set of GLUE benchmark. All the results reported here are on mini-1k. Bold values represent the best amongst synthetic splits
Fine-tuning on small amounts of Web Scraped Data boosts performance: Even after filtering, we observe that language models trained solely on synthetic text slightly underperform LMs trained on clean text. To address this issue, we conduct extended pretraining of LMs using clean data sourced from IndicMonoDoc. The objective is to determine if this additional training improves performance. We only incorporate an additional 10% of clean data compared to the LM’s previous training data. We see these results across all three languages, and for Hindi and Gujarati, we see that by incorporating even a small amount of clean data, we observe an increase in performance on downstream tasks, bringing the LM at par or closer to the performance of a clean LM. We see an improvement in LMs trained using unfiltered synthetic corpora as well but we believe that filtering leads to the removal of noisy documents and thus better performance. We observe improved performance in language models (LMs) trained with unfiltered synthetic corpora, but filtering out noisy documents enhances performance further. Our ablation study (Table 13 in Appendix A) investigates whether adding 10% more synthetic data contributes to this improvement or if the data type is key. While performance gains could stem from statistical variances, the consistency across nearly all downstream tasks suggests otherwise.
[…]
Model iXNLI bbc-a iitp-mr iitp-pr midas Avg.
HI-clean 68.74 80.25 67.74 77.05 78.33 74.42
syn-HI_en-unfiltered 67.32 77.92 65.63 76.81 77.58 73.05
syn-HI_en-filtered 69.48 78.98 65.16 77.43 77.33 73.68
syn-HI_en-filtered+1070.15 79.56 67.09 78.2 79.03 74.81
Table 3: Effect of reducing model size for Hindi on IndicGLUE accuracy. All the results reported here are on mini-1k. Bold values represent the best amongst synthetic splits
5.2 Further Exploration
Analysis of Synthetic Data: Figure LABEL:fig:_variance_plot shows the perplexity mean and variance scores for TinyLM across token positions in the documents. This shows that on unseen documents, TinyLM shows higher variance on English documents generated by translating Gujarati documents from IndicMonoDoc as compared to English clean and English synthetic generated from Hindi. This also gives reason for the deterioration in results in Table 2 due to Gujarati documents. Figure 7 shows the distribution of lengths of filtered documents by TinyLMs showing that they do not add any bias for shorter documents during filtering.
Impact of model size: Following Table 2 and 3, we see that even after scaling down we see consistent improvements for filtering and adding additional data, which empirically shows that indeed using synthetic text after filtering is a viable option for pretraining LMs of varying sizes. In Table 3 we see that after filtering and extended pretraining, synthetic text outperforms LMs trained on clean documents from the web in Hindi. This is also supported by our experiments on finetuning Llama-3-8B in Section 5.3.
Model XLSum HG XLSum QG Cnn Dialogsum Avg.
EN-clean 23.87 24.05 16.08 20.39 21.10
syn-EN_hi-unfiltered 22.17 22.97 12.56 18.30 19.00
syn-EN_hi-filtered 23.27 23.83 15.88 19.83 20.70
Table 4: Performance of English models on NLG tasks. All the results reported here are on base-1k and use Rouge-L F1 scores.
Impact on NLG: Without extended pretraining, language models trained on synthetic text perform as well as those trained on clean documents, suggesting that for NLG tasks, synthetic data suffices for pretraining, eliminating the need for clean data. This trend is evident across Hindi, Gujarati, and English NLG results (Tables LABEL:tab:hi-gu-main and 4). As their performance matches models trained on clean data, we refrain from extended pretraining for NLG tasks, focusing primarily on abstractive summarization for evaluating generation capabilities.
5.3 Scaling to Llama-3-8B
To show the effect of using translationese on larger models, we select Llama-3-8B [^12] https://github.com/meta-llama/llama3 and Gemma-2B (Team et al., 2024) and perform continual pretraining over clean and synthetic data to improve ability over the low resource target language. We take Marathi as a replacement for Hindi for scaling experiments since data for Hindi is abundantly available and existing models already have a good language understanding of Hindi making it harder to compare the effects of utilizing clean vs synthetic data. For a fair comparison, we limit each data split to 344M tokens for Gujarati and 465M tokens for Marathi and follow a similar procedure as described in Section 3.4 to generate and filter data for Marathi. We perform extended training for a single epoch using LoRA (Hu et al., 2021) finetuning on $W_{q}$, $W_{v}$ projection matrices using $\alpha$=$16$ and $r$=$8$. We keep the learning rate at $3e^{-5}$ with a weight decay of $0.01$ and an effective batch size of 58k.
Perplexity: We report average sentence level perplexity on sentences from IN22-Conv and IN22-Gen Gala et al. (2023) in Table 5. We see that filtered synthetic data for Gujarati outperforms clean data, but for Marathi, it does not. This means that filtering improves performance at scale but relies on the quality of translationese in the target language. We report perplexities on individual test sets in Appendix A.
Data Marathi Gujarati
Gemma
2B
Llama-3
8B
Gemma
2B
Llama-3
8B
Base model 178.898 66.740 71.136 2.839
clean 37.599 11.196 10.350 2.312
synthetic-unfiltered 92.813 15.697 10.941 2.816
synthetic-filtered 104.148 14.622 10.150 2.236
Table 5: Average perplexity ($\downarrow$) of models trained on Translationese vs. Clean data on IN22-Conv and IN22-Gen test sets shows improvement with large-scale models. Bold represents best among synthetic data splits.
[…]
Data Flores-200 IN22-Gen
Marathi Gujarati Marathi Gujarati
Base model 27.83 34.35 30 34.92
clean 34.02 35.61 33.94 35.42
synthetic-unfiltered 30.63 34.19 29.67 32.1
synthetic-filtered 31.81 35.54 31.3 35
Table 6: chrF++ scores on 5-shot Machine Translation FloRes and IN22-Gen test sets on Llama-3-8B. Bold represents best among synthetic data experiments.
Data Marathi Gujarati
Gemma
2B
Llama-3
8B
Gemma
2B
Llama-3
8B
Base model 90.89 ±0.005 95 ±0.009 83.84 ±0.002 92.69 ±0.018
clean 90 ±0.014 97.17 ±0.013 87.79 ±0.012 93.33 ±0.0132
synthetic
unfiltered
89.66 ±0.012 95.38 ±0.016 83.88 ±0.009 92.81 ±0.01
synthetic
filtered
86.67 ±0.016 96.15 ±0.011 84.10 ±0.013 92.94 ±0.007
Table 7: Average accuracy with standard deviation (superscript) over 5 runs on 10-shot IndicSentiment classification task. Bold represents the best among synthetic data experiments.
6 Conclusion
In this paper, we performed a first of its kind study showing the promise of using translationese data for training language models for low-resource languages. Our simple pipeline involves the translation of high-resource source language documents at scale, followed by perplexity based filtering using small and efficient language models trained on clean target low-resource language data. We then showed on a variety of downstream natural language understanding and generative tasks that both small and large language models pre-trained on clean synthetic data are comparable to those trained on clean data. While we observed that the source language for synthetic data generation matters, it is clear that synthetic data can help bridge the resource scarcity faced by a vast majority of languages for language modeling. Future work will focus on better and faster synthetic data generation and filtering mechanisms.
Acknowledgements
We extend our thanks to the Department of Computer Science and Engineering at the Indian Institute of Technology (IIT) Bombay for providing access to GPU servers, and to the Centre for Development of Advanced Computing (C-DAC) for granting us access to the Param Siddhi Supercomputer. These computational resources were essential for the successful completion of this work.
Limitations
We consider the following limitations of our work.
  • We show that synthetic data also helps for larger models like Llama-3-8B but for even larger models above 100B parameters, effects of translationese may be different. However, synthetic data generated from translations can surely help fill knowledge gaps.
  • Due to the extensive size of the test sets for IndicNLG tasks (Question Generation, WikiBio generation, Headline Generation, and Sentence Summarization), we couldn’t experiment with them in their entirety. However, since we already use 4000 examples per language, we anticipate that the overall trends remain unchanged.
  • We report GLUE validation set results for all models due to the large scale of our experiments, following existing practices. Our goal is to demonstrate synthetic data utility, not to achieve state-of-the-art results.
  • Our framework heavily relies on the translation model’s performance. Despite this dependency, we are confident that our approach will significantly enhance the performance of mid-resource languages, especially where the translation model is already proficient.
Ethical Considerations
As a part of this paper, we release monolingual and synthetic data. While we have taken care to remove any toxic content, accidental occurrences may exist and thus we exercise caution when using our data for training language models as they may produce toxic outputs. Given that we have shown the utility of synthetic data for training LMs, it should be possible to mass produce synthetic toxic data in various languages leading to LMs that can generate multilingual toxic content. However, this opens up research opportunities on how to detect and filter toxic content from synthetically created data.
We release the code and models with an MIT License [^13] https://opensource.org/license/mit/. The dataset is released under a CC-0 License [^14] https://creativecommons.org/share-your-work/public-domain/cc0/.
Appendix A Additional results
We report additional results in this section. Tables 14, 15 show the chrF++ and BLEU scores across three translation evaluation benchmarks. This shows that using parallel synthetic data does not deteriorate the performance of the language model. Similar results are shown in Table 16 for IndicNLG tasks where performance on Hindi generation tasks are only affected by a small margin and coupled with results in Table 8 showing that scores are not affected by using Hindi synthetic parallel data.
sst2 cola mrpc qnli qqp rte mnli-m mnli-mm stsb
Model acc mcc f1 acc f1 acc acc acc pearson Avg.
EN-clean 90.94 40.26 87.4 84.98 84.47 65.34 77.84 77.96 82.67 76.87
syn-EN_hi-unfiltered 84.61 31.1 81.78 79.35 81.44 63.3 72.94 73.16 78.9 71.84
syn-EN_hi-unfiltered + 10% 87.39 34.22 85.77 80.96 81.07 65.11 74.76 74.38 80.32 73.78
syn-EN_hi-filtered 88.3 34.03 86.55 83.59 83.64 63.17 75.6 75.41 81.1 74.60
syn-EN_hi-filtered + 10% 90.13 35.75 86.41 84.75 84.21 65.34 76.99 76.91 81.95 75.83
Table 8: Results on English: Dev set of GLUE tasks for different synthetic splits on the base-1k model. Synthetic LMs perform almost as well as clean LMs after filtering and further training with clean data. Bold values represent the best amongst synthetic splits.
Using synthetic for one language doesn’t impact performance in another: For many multilingual language models, data imbalance causes a gap in performance across languages. But what if we can combine synthetic data along with clean data for training multilingual models? would the synthetic part deteriorate the performance of the multilingual model? To experiment with this, we train bilingual base-1k models over different combinations of clean and synthetic corpora for English and Hindi and evaluate their performance on GLUE Wang et al. (2018), and report performance on IndicNLG, and Machine translation in Appendix A. Following Table 9, we see that using Hindi synthetic data does not affect its performance compared to BI-EN-HI-clean model which is solely trained on clean corpora. This implies that it is possible to train multilingual models where some languages are trained only over a clean subset and others on synthetic without deteriorating performance across languages. We further see that using parallel data does not have much impact on multilingual models.
sst2 cola mrpc qnli qqp rte mnli-m mnli-mm stsb
Model acc mcc f1 acc f1 acc acc acc pearson Avg.
BI-EN-HI-clean 89.56 38.53 85.56 84.88 84.39 64.25 76.4 77.27 82.07 75.88
BI-EN-HI_syn-parallel-filtered 89.56 39.57 85.71 84.75 84.62 64.98 77.31 77.85 82.41 76.31
BI-EN-HI_syn-nonparallel-filtered 89.79 38.68 86.92 85.08 84.06 65.34 77.15 77.55 83.01 76.40
BI-EN_syn-HI_syn-filtered 87.95 30.05 84.9 83.7 83.97 63.89 75.63 76.24 82.24 74.29
BI-EN_syn-HI_syn-filtered + 10% 89.1 35.45 85.34 84.53 84.18 65.7 76.64 77.24 82.1 75.59
Table 9: Results on English for Bilingual models: Dev set of GLUE tasks for different synthetic splits on the base-1k model. Training bilingual models using synthetic data in one language (Hindi) does not affect the performance in the other language (English). Bold values represent the best amongst synthetic splits.
FLORES
Model EN-HI HI-EN Avg.
BI-EN-HI-clean 46.56 51.7 49.13
BI-EN-HI_syn-parallel-filtered 44.12 50.64 47.38
BI-EN-HI_syn-nonparallel-filtered 45.65 51.29 48.47
EN-GU GU-EN Avg.
BI-EN-GU-clean 26.44 35.3 30.87
BI-EN-GU_syn-parallel-filtered 26.77 34.84 30.81
BI-EN-GU_syn-nonparallel-filtered 26.7 36.54 31.62
Table 10: chrF++ Scores on FLoRes translation task. EN-HI models are based on base-1k and EN-GU models are based on mini-1k
Impact on Machine Translation: (MT) We focus on MT separately as a special case of NLG. We hypothesized that using parallel synthetic documents for bilingual models would improve translation performance by enhancing alignment between languages. However, our evaluation fails this hypothesis. Results indicate that using nonparallel synthetic documents yields similar translation performance across language directions and benchmarks compared to parallel synthetic documents. This might be because there is no explicit alignment happening during training between parallel documents. See Table 10 for chrF++ scores on FLoRes-200 Team et al. (2022), and Appendix A for chrF++ and BLEU scores on IN22-Conv, IN22-Gen Gala et al. (2023).
Perplexity-IN22 Conv
Data Marathi Gujarati
Gemma-2B Llama-3-8B Gemma-2B Llama-3-8B
Base model 332.9036 121.6586 131.8027 3.3922
clean 58.3804 15.1342 12.3128 2.5208
synthetic-unfiltered 168.0933 23.018 12.7995 3.0671
synthetic-filtered 191.1506 21.355 11.6916 2.3849
Table 11: Perplexity on IN-22 Conv. Bold values represent best among synthetic splits.
Perplexity-IN22 Gen
Data Marathi Gujarati
Gemma-2B Llama-3-8B Gemma-2B Llama-3-8B
Base model 24.893 11.821 10.4693 2.2853
clean 16.8172 7.2572 8.3868 2.1032
synthetic-unfiltered 17.5333 8.3766 9.0821 2.564
synthetic-filtered 17.1451 7.8885 8.6089 2.0868
Table 12: Perplexity on IN-22 Gen. Bold values represent best among synthetic splits.
NLU NLG
Model iXNLI bbc-a iitp-mr iitp-pr midas Avg.
Headline
Gen.
Sentence
Summ.
Question
Gen.
Wikibio Avg.
HI-clean 73.61 81.75 72.58 79.73 80.34 77.60 27.54 23.64 24.84 52.16 32.04
syn-HI_en-unfiltered 72.87 77.92 64.36 76.22 79.91 74.26 27.29 22.93 24.22 50.14 31.14
syn-HI_en-unfiltered+10clean 74.63 78.36 67.75 77.46 80.17 75.67 26.98 23.20 24.76 51.34 31.57
syn-HI_en-filtered 74.75 81.06 69.03 78.58 79.73 76.63 27.15 23.10 24.41 49.88 31.13
syn-HI_en-filtered+10clean 74.49 80.94 71.61 79.92 80.64 77.52 27.87 24.23 24.87 51.18 32.04
syn-HI_en-filtered+10synthetic 74.95 80.83 71.93 79.92 78.73 77.27 27.64 25.13 23.60 43.48 29.96
Table 13: Ablation Results for Hindi using additional 10% clean vs synthetic data: NLU/NLG tasks on base-1k (Hindi) different clean and synthetic splits. Test accuracy for NLU tasks; Rouge-L F1 scores for NLG tasks. Bold values represent the best amongst synthetic splits.
IN22-Conv IN22-Gen FLORES
Model EN-HI HI-EN EN-HI HI-EN EN-HI HI-EN
BI-EN-HI-clean 41.22 50.3 43.49 47.83 46.56 51.7
BI-EN-HI_syn-parallel-filtered 41.92 49.67 41.61 46.95 44.12 50.64
BI-EN-HI_syn-nonparallel-filtered 40.74 49.54 42.28 47.66 45.65 51.29
EN-GU GU-EN EN-GU GU-EN EN-GU GU-EN
BI-EN-GU-clean 35.85 41.27 22.95 31.83 26.44 35.3
BI-EN-GU_syn-parallel-filtered 34.36 41.86 22.93 30.84 26.77 34.84
BI-EN-GU_syn-nonparallel-filtered 34.49 42.08 23.06 32.81 26.7 36.54
Table 14: chrF++ Scores on FloRes, IN22-Conv and IN22-Gen splits for translation task. EN-HI models are based on base-1k and EN-GU models are based on mini-1k. Bold values represent the best amongst synthetic splits.
IN22-Conv IN22-Gen FLORES
Model EN-HI HI-EN EN-HI HI-EN EN-HI HI-EN
BI-EN-HI-clean 19.58 23.01 17.23 19.72 21.8 21.73
BI-EN-HI_syn-parallel-filtered 19.64 23.79 16.57 20.14 21.63 22.6
BI-EN-HI_syn-nonparallel-filtered 19.25 22.47 16.37 19.74 21.51 21.74
EN-GU GU-EN EN-GU GU-EN EN-GU GU-EN
BI-EN-GU-clean 10.24 15.19 4.65 7.92 5.44 9.57
BI-EN-GU_syn-parallel-filtered 11.24 15.7 4.87 8.44 6.7 10.02
BI-EN-GU_syn-nonparallel-filtered 10.86 15.57 5.07 9.07 6.17 10.03
Table 15: BLEU Scores on FloRes, IN22-Conv and IN22-Gen splits for translation task. EN-HI models are based on base-1k and EN-GU models are based on mini-1k. Bold values represent the best amongst synthetic splits.
Model Headline Generation Sentence Summarization
Question
Generation
Wikibio
Generation
BI-EN-HI-clean 27.47 23.78 24.25 50.82
BI-EN-HI_syn-parallel-filtered 26.96 23.10 25.38 48.26
BI-EN-HI_syn-nonparallel-filtered 27.32 22.84 24.95 50.22
Table 16: Performance of Bilingual models on IndicNLG tasks. All the results reported here are on base-1k and use Rouge-L F1 scores. Bold values represent the best amongst synthetic splits.
IndicXNLI 10-Shot
Data Marathi Gujarati
Gemma-2B Llama-3-8B Gemma-2B Llama-3-8B
Base model 36.56 ± 0.47 50.66 ± 1.46 34.74 ± 0.83 47.63 ± 1.96
clean 35.92 ± 0.79 48.47 ± 1.44 33.40 ± 0.29 48.01 ± 0.99
synthetic-unfiltered 35.66 ± 0.93 49.50 ± 1.76 33.26 ± 0.39 48.22 ± 1.88
synthetic-filtered 34.4 ± 0.23 51.98 ± 1.32 33.78 ± 0.22 47.41 ± 2.17
Table 17: Classification Results on IndicXNLI using 10-shot prompting.
Appendix B Training and Evaluation
In this section, we provide an overview of the training and evaluation setup employed in our experiments. This includes details about the datasets used, training hyperparameters, evaluation metrics, and other relevant configurations.
B.1 Evaluation Datasets
For evaluation, we utilize a diverse set of datasets covering four languages: English, Hindi, Gujarati, and Marathi. For Hindi and Gujarati, we rely on the IndicGLUE benchmark [^15] https://huggingface.co/datasets/ai4bharat/indic_glue Kakwani et al. (2020b), which provides a range of tasks for natural language understanding (NLU), including natural language inference (IndicXNLI/iXNLI), article genre classification (bbc-a, iNLTK), discourse mode classification (MIDAS), and sentiment analysis (iitp-mr, iitp-pr). For natural language generation (NLG), we employ the IndicNLG benchmark [^16] https://huggingface.co/collections/ai4bharat/indicnlg-66c5a1397bab135be074cfe1 Kumar et al. (2022), which includes tasks like headline generation, sentence summarization, question generation, and Wikipedia biography generation.
The IndicGLUE dataset is semi-automatically curated using website metadata and Wikipedia articles, while IndicNLG is derived from Wikipedia articles and news websites for summarization, along with parallel corpora and pivot-based translation for paraphrasing tasks. Additionally, we incorporate the test sets from IN22 Gala et al. (2023) and Flores-200 Team et al. (2022) to evaluate performance on machine translation tasks.
We use the well-known GLUE benchmark, which includes nine NLU tasks in English such as natural language inference (NLI), semantic similarity, text classification, and linguistic acceptability. For English summarization tasks, we rely on XLSum Hasan et al. (2021), DialogSum Chen et al. (2021), and CNN/DailyMail See et al. (2017).
B.2 Training
For the pretraining of the base models, we keep a hard limit for the base-1k model as 2.38B tokens and for the mini-1k model as 1B tokens. But for the TinyLM we relax this token limit until we see overfitting. For our experiments, we use the NVIDIA A100-SXM4-80GB GPUs.
Hyperparameter Value
vocab_size 56000
val_every 0.05
bs 48
n_embed 768
num_blocks 4
num_heads 16
head_size n_embed // num_heads
context_len 1024
block_size context_len
attn_drop_value 0.1
dropout 0.1
ffn_drop_value 0.1
use_flashattn TRUE
ffn_scaling 4
positional_embedding rope’
rotatory_embedding_dim head_size // 2
lr 6.00E-04
wd 1.00E-01
beta_1 0.9
beta_2 0.95
eps 1.00E-05
epochs 2
precision bf16
accumulate_grad_batches 8
gradient_clip_val 1
strategy ddp’
accelerator gpu’
warmup_steps 5000
num_workers 16
SHUFFLE_SEED 42
PIN_MEMORY TRUE
NUM__NODES 1
NUM_DEVICES 2
Table 18: Hyperparameters used for training the mini-1k model
Hyperparameter Value
vocab_size 56000
val_every 0.05
bs 48
n_embed 768
num_blocks 12
num_heads 12
head_size n_embed // num_heads
context_len 1024
block_size context_len
attn_drop_value 0.1
dropout 0.1
ffn_drop_value 0.1
use_flashattn TRUE
ffn_scaling 4
positional_embedding rope’
rotatory_embedding_dim head_size // 2
lr 6.00E-04
wd 1.00E-01
beta_1 0.9
beta_2 0.95
eps 1.00E-05
epochs 2
precision bf16
accumulate_grad_batches 8
gradient_clip_val 1
strategy ddp’
accelerator gpu’
warmup_steps 5000
num_workers 16
SHUFFLE_SEED 42
PIN_MEMORY TRUE
NUM__NODES 1
NUM_DEVICES 2
Table 19: Hyperparameters used for training the base-1k model
B.3 Extended pretraining
For the mini-1k models, we randomly sample 100M tokens from the clean subset of IndicMonoDoc for the target language, and for the base-1k model, we sample 200M for extended pretraining. We use the same hyperparameters as training and perform extended pretraining for 2 epochs over this newly sampled clean data. For scaling experiments, we utilize TorchTune [^17] https://github.com/pytorch/torchtune for fine-tuning Llama-3-8B and Gemma-2B models. For a fair comparison, we limit each data split to 344M tokens for Gujarati and 465M tokens for Marathi and follow a similar procedure as described in Section 3.4 to generate and filter data for Marathi. We perform extended training for a single epoch using LoRA (Hu et al., 2021) finetuning on $W_{q}$, $W_{v}$ projection matrices using $\alpha$=$16$ and $r$=$8$. We keep the learning rate at $3e^{-5}$ with a weight decay of $0.01$ and an effective batch size of 58k. We use the AdamW optimizer Loshchilov and Hutter (2019) with 1000 warmup steps and a cosine learning rate scheduler.
B.4 Fine-tuning
For GLUE tasks we use the dev split on the clean part and do hyperparameter tuning to achieve the best scores, and then we use the same hyperparameters for all synthetic experiments. For IndicGLUE we follow a similar setting for the val split to find good hyperparameters and report results on the test split like Kakwani et al. (2020a). For all classification and regression tasks, we use a single linear layer and use an appropriate activation function for classification and regression respectively. We use an Adam optimizer Kingma and Ba (2015) with a learning rate of $1e^{-5}$ and a batch size of 48. For NLG tasks we do extended pretraining using a separator token in between the input and output sequence with an effective batch size of 768 examples and only calculate loss for the output sequence. We use an AdamW optimizer Loshchilov and Hutter (2019) with learning rate = $6e^{-4}$, weight decay = $1e^{-1}$, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95 and $\epsilon$ = $1e^{-5}$. For translation, we randomly sample 1M parallel sentence for each language pair from the samanantar corpus Ramesh et al. (2022) and evaluate on FloRes Team et al. (2022), IN22-Conv and IN22-Gen Gala et al. (2023). We list the batch size and number of epochs of each task in Table 20.
B.5 Prompting
We use random sampling from validation sets whenever available and utilize other examples from the test set otherwise. We take 5 random samples for each evaluation of NLG tasks and 10 random samples for each evaluation of NLU tasks. We list down the prompt used below for Marathi evaluations, we use similar prompts for Gujarati as well.
[…]
Task Batch size Epochs Metric
IndicXNLI 48 5 Accuracy
BBC-Articles 24 20 Accuracy
IITP-MR 24 20 Accuracy
IITP-PR 48 20 Accuracy
MIDAS 48 20 Accuracy
Headline Generation 768 2 Rouge-L F1
Sentence Summarization 768 2 Rouge-L F1
Question Generation 768 2 Rouge-L F1
WikiBio Generation 768 4 Rouge-L F1
iNLTK 48 20 Accuracy
sst2 48 10 Accuracy
CoLA 48 30 MCC
mrpc 48 30 F1
qnli 48 10 Accuracy
qqp 48 5 F1
rte 48 30 Accuracy
mnli-matched 48 5 Accuracy
mnli-mismatched 48 5 Accuracy
stsb 48 20 Pearson
XLSum Headline Gen. 768 4 Rouge-L F1
XLSum Question Gen. 768 4 Rouge-L F1
CNN Dailymail 768 4 Rouge-L F1
DialogSum 768 4 Rouge-L F1
Samanantar 768 2 chrF++ / BLEU
Table 20: Hyperparameters used for finetuning tasks
B.6 Evaluation
We use torch metrics [^18] https://lightning.ai/docs/torchmetrics/stable/pages/lightning.html to calculate accuracy, f1-score, Pearson correlation, Matthew’s correlation coefficient. We report chrF++ scores191919chrF++ signature
nrefs:1—case:mixed—eff:yes—nc:6—nw:2—space:no—version:2.4.0
and BLEU scores202020sacreBLEU signature:
nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.4.0
Papineni et al. (2002) using the sacreBLEU [^21] https://github.com/mjpost/sacrebleu implementation and Rouge-L f1 scores using the sacreRouge Deutsch and Roth (2020) implementation by the xl-sum repository [^22] https://github.com/csebuetnlp/xl-sum.
We report English scores for NLU on the validation split of the GLUE benchmark and test splits for XL-Sum, CNN Dailymail, and Dialogsum NLG benchmarks. For Hindi and Gujarati, we use the test split of IndicGLUE and IndicXNLI.
For classification and regression tasks, we use the models finetuned according to hyperparameters mentioned in Appendix B.4 to keep fair comparison for all models and mention results on the final epoch. For generations on IndicNLG and English NLG tasks, we use beam search with a beam width of 5, length penalty of 1.0, n_gram repetition penalty of 4 n_grams with sampling set to false and early stopping set to true. We also set a maximum generation length to 64 tokens. For the translation task, we follow a beam search with a beam width of 5, maximum new tokens to 256 and early stopping to true.
Appendix C Utilising Translationese
In this section we provide details on the generation and filtering of translationese data for our experiments.
C.1 Creating synthetic data
“Translationese” is a term used to describe peculiarities in the text translated into a specific language, differentiating it from content originally written in that language Gellerstam (1986). Translated texts into the target language (via humans or machine-generated) often show distinctive features that differentiate them from their original counterparts in the target language. These disparities arise from either the influence of the translation process itself on the final product or the inherent “fingerprints” of the source language subtly present in the target language rendition Rabinovich and Wintner (2015). This is a common phenomenon in translation models where the target language translations often show characteristics of the source language and add bias to the evaluation of downstream tasks Toral et al. (2018); Zhang and Toral (2019); Graham et al. (2019). So far a lot of work on synthetic translated data has been done for using back translations Sennrich et al. (2016a); Edunov et al. (2018) for improving Machine translation performance Marie et al. (2020); Bogoychev and Sennrich (2019); Ni et al. (2022) or for classification tasks like native language identification Goldin et al. (2018), etc. Tranlationese data has been used for many tasks but we explore the efficacy of using translationese data for pretraining of language models. We collect monolingual corpora in the source language as mentioned in Section 3.1 and utilize a powerful off-the-shelf translation model IndicTrans2 Gala et al. (2023) to generate translationese data. Since IndicTrans2 can only handle a max sentence length of 256 BPE tokens, we split the documents using Moses Sentence Splitter [^23] https://pypi.org/project/mosestokenizer/ to perform translations into the target language at the sentence level and then merge again to form documents. We also repair translations that exceed in length 256 BPE tokens using the TinyLM trained on clean corpora as mentioned in Section 4 to complete the sentence translation, we encounter only 0.002% of such cases. We use this corpus for the synthetic and clean+synthetic part of our experiments.
Figure 7: Violin plot displaying the distribution of lengths of clean and filtered English documents on different data splits: en-clean (English web documents), syn-en_hi (synthetic English documents translated from Hindi), and syn-en_gu (synthetic English documents translated from Gujarati).
C.2 Perplexity filtering
Following Figure 2, we use these TinyLMs to filter the generated synthetic translationese corpora from IndicTrans2. We do this by using perplexity as a measure of document quality score. For language models, perplexity quantifies how well a model predicts a sequence of tokens. A lower perplexity indicates better predictive performance. While calculating perplexity over a sequence of tokens $W\in$ $w_{1},w_{2},\ldots,w_{N}$ we skip the first $s$ tokens where $s=10$, $e=1024$ and calculate loss until only the first $e$ tokens of the document. We find setting $e$ to larger values can lead to higher variance in the document scores due to the size of the TinyLM. After initial analysis, we choose $s$ and $e$ such that we remove the high uncertainty of the language at the start of an unseen document and avoid penalizing longer documents due to the fragility of the extrapolation ability of TinyLM [^24] During experiments we saw that these TinyLMs can only go up to a certain context length before deteriorating in quality.. Note that it is important to choose $e$ such that the language model gives a uniform estimate of perplexity over an already seen sequence of tokens $\in$ $w_{s},w_{s+1},\ldots,w_{e}$. For our experiments, we use the TinyLMs to score all synthetically generated translationese data and calculate a document score using the above method. Following Laurençon et al. (2022), we do subsampling by thresholding document perplexity scores except Laurençon et al. (2022) did it using Ken-LM Heafield (2011) and we do it using our TinyLM. We keep the threshold value such that we include enough documents to reach the computed optimal token count for pretraining experiments.
High Resource Low Resource
Lang #Tokens Lang #Tokens
bn 5,258.47 as 57.64
en 11,986.53 brx 2.25
gu 887.18 doi 0.37
hi 11,268.33 gom 2.91
kn 567.16 kas 1.27
ml 845.32 mai 1.51
mr 1,066.76 mni 0.99
ne 1,542.39 or 81.96
pa 449.61 sa 80.09
ta 2,171.92 sat 3.05
te 767.18 sd 83.81
ur 2,391.79
Table 21: Languagewise corpora size in Million tokens
Appendix D Qualitative Analysis
Since translation errors occur frequently, leading to biased, ungrammatical, or erroneous translations that can have drastic consequences on training, it is important to mitigate or remove such errors in translationese corpora. The most common machine translation errors include mistranslations due to ambiguous words, incorrect handling of expressions, syntax and grammar errors, and issues with preserving context across longer sentences. Many approaches have been proposed to address these errors, but most are computationally expensive, especially when translationese data is used for pretraining. Instead, we ask whether a language model can identify such issues. To investigate this, we examine which types of English sentences were filtered.
In many cases, we found that the filtered documents included errors like code-mixing and repetitions, often generated by the diverging output of the translation model. This is expected since such phenomena are rarely seen in natural written language, and the model assigning high entropy suggests an unlikely sequence outcome. Although the model regarded some erroneous instances as false positives, many such cases were successfully avoided. We also noticed that, due to the small size of the model, many perfectly good documents were filtered out because of the model’s inability to evaluate them. This issue was observed less frequently in larger models used for evaluation. As seen in Figure 8, the filtering model fails to understand complex words and named entities, which it regarded as unlikely due to its preference for simpler terms and more likely entities. Other common elements that were filtered included numbers, dates, and abbreviations. While this can lead to the loss of valuable information, as many good documents are discarded, we empirically observe the benefits of such filtering. These errors could likely be reduced by using a larger filtering model that can better approximate the source language but we leave this analysis for future work.
Figure 8: Heatmap of perplexity over filtered sentences.
Appendix E IndicMonoDoc
In this section, we describe the process of creating the IndicMonoDoc corpus which is the largest document-level corpora for Indic languages consisting of 39.5 billion tokens spanning 23 languages. IndicMonoDoc comprises 27.5B Indic tokens and 12B tokens of English tokens. Table 21 shows language language-wise deduplicated size of the IndicMonoDoc corpus and Figure 3 shows a comparative 100% stacked bar plot with IndicCorpv2 which is a sentence level corpora.
E.1 Crawling
To extract URLs from the web we sample word level n-grams; n=2,…,6 from a sample monolingual corpora to create a list of keyword searches. We then randomly merge k; k=1,..,4 keywords to form a query. Using these queries we perform automatic web searches to collect a large repository of URLs. We merge this list with a manual list of sources to perform URL-level deduplication. We crawl these webpages leaving out some of them [^25] We leave webpages consisting of a robots.txt file and URLs containing offensive text or social media links. We leave out webpages that consist of a considerable amount of English content using a simple script recognition regex. We perform this scrapping majorly for the bottom 14 low-resource languages. We also add script-level recognition using Unicode characters [^26] https://unicode.org/charts/ for each language before crawling a webpage to avoid scrapping non-Indic text.
E.2 Post processing
A lot of crawled content consists of unwanted text like HTML tags, emoticons, and text in another language. We use manual filtering pipelines inspired by OSCAR Ortiz Suárez et al. (2019), Abadji et al. (2022) to remove such content. We additionally use a language detection-based (LID) filtering using cld3272727https://github.com/google/cld3 and IndicLID-FTN model Madhani et al. (2023a) to discard languages not of interest. Following Doddapaneni et al. (2023) we perform document filtering to remove offensive text from the corpora using a list of offensive words and phrases extended from work by Team et al. (2022) which consists of offensive words in 209 languages. We also use a Romanized version of this list using the transliteration tool by Madhani et al. (2023b) to perform toxic document filtering in 17 languages. Following Kakwani et al. (2020a) & Doddapaneni et al. (2023) we merge all the filtered corpus with Wikipedia, OSCAR Ortiz Suárez et al. (2019) and some dumps of mC4 Xue et al. (2021). Finally, we perform deduplication at paragraph level using Murmurhash algorithm [^28] https://pypi.org/project/mmh3/ with a 128-bit unsigned hash for each monolingual split of the corpora. After all post-processing steps, the language wise size of the corpora is mentioned in Table 21. A major chunk of the corpus is comprised of English, Hindi, and Bengali which make up 72.15% of the corpora.